# Bob’s Confetti: Phonetic Memorization Attacks in Music and Video Generation

### v3

@@ 4.4 AVT Attack Results
Table 4: AudioJudge and MiRA similarity scores for lyric-based song recreations. Melody, Rhythm, Harmony, Identity, and Style are from AudioJudge (gpt-4o-audio-preview); CLAP and CoverID are from MiRA. English Billboard songs were generated with YuE, Cantonese songs with SUNO. Genre prompts: Basket Case – none; Thinking Out Loud – “male romantic vocal guitar ballad with piano melody”; Let It Be, Billie Jean, Empire State of Mind, Lose Yourself – “inspiring female uplifting pop airy vocal electronic bright vocal vocal”; Cantonese songs – ballad-style prompt.
@@
To examine whether these verbatim prompts also trigger memorized behavior in the multimodal setting, we apply the AVT attack to Veo 3. When prompted with the exact lyrics of Lose Yourself, the model produced an even closer visual reproduction than under APT: a male rapper in a hoodie, placed in dimly lit, urban settings—closely mirroring the original music video’s aesthetic (Figure 4 of Appendix B). Notably, the tone, voice, and rhythm of the generated audio were also strikingly aligned with the original track, further reinforcing the presence of multimodal memorization. Similarly, for Jingle Bells Veo 3 consistently generated music that was melodically and rhythmically faithful to the original. This highlights the model’s strong tendency to regurgitate memorized content when exposed to exact training examples, extending lyric-based memorization across both audio and visual outputs.
4.5 Ablation Study
To better understand the mechanisms underlying memorization in L2S and T2V models, we conduct a series of controlled ablation studies that isolate the effects of phonetic similarity versus verbatim content. By varying genre prompts, lyric fidelity, and phonetic perturbations across matched and mismatched inputs, we aim to disentangle the respective contributions of surface form, semantic content, and phonetic structure in triggering memorized generations. These studies expose the robustness and modality-transferability of memorization behaviors in modern generative models.
Genre Prompt Variation.
Even without any stylistic conditioning, YuE reproduces audio that closely aligns with training data when the lyrics match known examples. For instance, in the Mandarin song 天后 (Andrew Tan), AudioJudge assigns strong scores (melody = 0.88, rhythm = 0.85), while MiRA reports CLAP = 0.638 and CoverID = 0.300, indicating overlap with memorized content. This pattern mirrors MiRA’s earlier observations of lyric-driven leakage, with AudioJudge now confirming that acoustic structure is also faithfully preserved under verbatim prompting. In addition, supplying the correct genre tag amplifies memorization. For example, 光辉岁月 (Beyond) retains high melodic and rhythmic fidelity (0.95 / 0.90), with MiRA reporting CLAP = 0.731 and CoverID = 0.401. Likewise, 海阔天空 (Beyond) achieves nearly identical scores (melody = 0.95, rhythm = 0.92, CLAP = 0.767), showing that genre alignment neither reduces nor meaningfully alters memorized outputs when lyrics remain unchanged (Table 7 of Appendix C.2).
Same Song, Different Genre.
To test whether YuE responds more strongly to stylistic prompts or lyric memorization, we generated 后来 (Rene Liu) under four genre conditions (Table 8). Despite prompts ranging from inspiring pop to gentle piano ballad, AudioJudge and MiRA scores remain tightly clustered (melody = 0.90–0.95, rhythm = 0.75–0.92, CLAP = 0.785–0.858, CoverID = 0.291–0.570). These stable results indicate that genre conditioning has limited influence over musical structure, with YuE’s generations overwhelmingly anchored to the lyrics themselves, further suggesting a strong lyric-driven overfitting to training data (Table 8 of Appendix C.2).
@@ 5 Discussion
Table 5: Cosine similarity between original and modified lyrics across YuE/GPT embeddings.
Song Embedding Cosine Sim.
Lose Yourself YuE 0.976
GPT 0.746
DNA YuE 0.960
GPT 0.725
APT YuE 0.513
GPT 0.755
@@
This interpretation is supported by embedding analyses (Table 5), which show that phoneme-preserving variants of rap tracks remain highly similar in YuE embeddings, reflecting the model’s reliance on rhythmic–phonetic alignment over semantic content. In contrast, genres where melody is the primary driver, such as modern K-Pop, showed lower sensitivity: phoneme-based prompts did not reproduce memorized outputs, despite the models’ strong performance in generating these songs.
@@
6 Conclusion
In this work, we introduce Adversarial PhoneTic Prompting (APT) attack, which exposes a new memorization vulnerability in L2S and T2V generation models. By altering lyrics to preserve phonetic structure while discarding semantics, we show that models like SUNO, YuE, and Veo 3 can reproduce memorized musical and visual content with high fidelity. These results highlight the model’s sensitivity to sub-lexical rhythm and cadence, revealing that phonetic cues alone—particularly in rhythmically structured genres like rap and iconic pop music—can serve as implicit triggers for memorization without lexical overlap or explicit cues. These findings expose an emerging risk in text-to-audio and transcript-conditioned generation pipelines, where phonetic form acts as a latent key to stored content. The success of our attack suggests that the demonstrated memorization behavior may emerge in transcript-conditioned generative systems, and we leave further investigation in this space of multi-modal generation for future work. As these models continue to be deployed in commercial and creative workflows, our results underscore the urgent need for new evaluation and safety frameworks that account for phonetic, rhythmic, and multimodal leakage paths, not just semantic or token-based similarity.
Appendix
Appendix A AudioJudge Heatmap
Figure 3: AudioJudge Similarity Heatmaps. We evaluate pairwise melody and rhythm similarity between original and generated songs using AudioJudge across four categories: (1) Mandarin, (2) Cantonese, and (3) other English songs. Each heatmap cell shows the overall similarity score (0–100) between an original and generated song. Green indicates high similarity (80–100), yellow moderate (40–79), and red low similarity (0–39). Diagonal cells reflect self-pairing scores (i.e., original with phoneme-modified versions of the same song). The distribution of scores confirms that AudioJudge does not assign uniformly high scores across all comparisons, but rather discriminates meaningfully based on melodic and rhythmic correspondence. This supports its reliability as an evaluative tool for music generation similarity.
Figure 3 provides a comprehensive visualization of AudioJudge’s similarity assessments across a diverse set of original and generated songs in four different evaluation scenarios: phoneme-modified English (top-left), Mandarin (top-right), English (bottom-left), and Cantonese (bottom-right). Each heatmap cell reflects the overall similarity score (range: 0–100), which aggregates melody and rhythm similarity scores produced by the GPT-4o-audio model given the AudioJudge prompt. To interpret the heatmaps: (i) Green cells (80–100) represent high similarity, (ii) Yellow cells (40–79) moderate similarity, and (iii) Red cells (0–39) low similarity.
Diagonal entries generally indicate the score between an original song and its own variant (e.g., a phoneme-modified or language-perturbed version). These diagonal scores are expected to be higher if the generation retains musical structure despite perturbations. Notably, the heatmaps demonstrate that AudioJudge does not trivially assign high similarity scores to all comparisons. For example, in the phoneme-modified group, "Let It Be" → phoneme variant receives a high similarity score (88), while unrelated pairs like "Can’t Help Falling" → "Lose Yourself" yield much lower scores (12–25). In the multilingual subsets (Mandarin and Cantonese), diagonal blocks exhibit high fidelity (e.g., “Houlai” → “Houlai”: 96), while cross-song scores drop significantly, reinforcing AudioJudge’s discriminative capability across tonal and rhythmic structure. The English subset further supports this, where "Basket Case" variants score 92 on the diagonal, yet cross-comparisons like "Empire State" → "Lose Yourself" yield much lower similarity (12–18). These patterns confirm that AudioJudge is sensitive to fine-grained audio alignment and does not exhibit mode collapse or over-averaging. This validates its use as a core similarity metric for identifying memorization phenomena in generated music.
Appendix B APT & AVT attack on Veo 3 Generation Example
(a) "Lose Yourself" Original MV Video Frame
(b) "Lose Yourself" Veo 3 Generation
(c) "Jingle Bell" Veo 3 Generation
Appendix C Additional Experimental Results
C.1 APT Attack: Christmas Songs Results

### v4

@@ 5.5 AVT Attack Results
Table 4: AVT Attack Results. AudioJudge and MiRA similarity scores for lyric-based song recreations. Melody, Rhythm, Harmony, Identity, and Style are from AudioJudge (gpt-4o-audio-preview); CLAP and CoverID are from MiRA. English Billboard songs were generated with YuE, Cantonese songs with SUNO. Genre prompts: Basket Case – none; Thinking Out Loud – “male romantic vocal guitar ballad with piano melody”; Let It Be, Billie Jean, Empire State of Mind, Lose Yourself – “inspiring female uplifting pop airy vocal electronic bright vocal vocal”; Cantonese songs – ballad-style prompt.
@@
To examine whether these verbatim prompts also trigger memorized behavior in the multimodal setting, we apply the AVT attack to Veo 3. When prompted with the exact lyrics of Lose Yourself, the model produced an even closer visual reproduction than under APT: a male rapper in a hoodie, placed in dimly lit, urban settings—closely mirroring the original music video’s aesthetic (Figure 4). Notably, the tone, voice, and rhythm of the generated audio were also strikingly aligned with the original track, further reinforcing the presence of multimodal memorization. Similarly, for Jingle Bells Veo 3 consistently generated music that was melodically and rhythmically faithful to the original. This highlights the model’s strong tendency to regurgitate memorized content when exposed to exact training examples, extending lyric-based memorization across both audio and visual outputs. Additional ablation studies examining genre prompt variation and cross-genre generalization are provided in Appendix H.
@@ 6 Discussion
6.1 AudioJudge Validation
To validate AudioJudge’s discriminative capability, we visualize similarity scores across original-generated song pairs for both APT and AVT attacks (Figure 5 and Figure 10 in Appendix D). Each heatmap cell represents the overall similarity score (0–100), aggregating melody and rhythm assessments from gpt-4o-audio-preview. The results confirm that AudioJudge reliably distinguishes memorized content from unrelated generations. Diagonal entries (original vs. its generated variant) show high similarity—65–92 for APT and 88–96 for AVT—while off-diagonal entries (cross-song comparisons) remain low (12–30). This separation demonstrates that AudioJudge does not trivially assign high scores to all comparisons, validating its use as a memorization metric. Notably, AVT consistently achieves higher diagonal scores than APT, confirming that verbatim lyrics trigger stronger memorization than semantically-altered prompts.
Figure 5: AudioJudge heatmap for APT attack. Diagonal entries show similarity between originals and their APT variants. High diagonal scores (65–92) demonstrate successful memorization triggering despite semantic alteration of lyrics.
6.2 Why APT Attacks Trigger Memorization
@@
Embedding Analysis.
To quantify the extent to which YuE’s text encoder captures phonetic versus semantic features, we compute cosine similarity between original lyrics and their APT-attacked lyrics across two embedding spaces: (1) YuE’s internal text embeddings, extracted via mean-pooling over the model’s input embedding layer, and (2) Sentence-BERT [reimers2019sentencebertsentenceembeddingsusing] embeddings (all-mpnet-base-v2), which capture semantic similarity. Figure 6 plots these similarities for 13 songs spanning multiple genres and languages. The results reveal a striking divergence: YuE embeddings show consistently high similarity ($\mu=0.90$, $\sigma=0.13$) between original and modified lyrics, while semantic embeddings show substantially lower similarity ($\mu=0.71$, $\sigma=0.19$). The low correlation between the two measures ($r=0.13$) confirms that they capture fundamentally different textual properties. This pattern is particularly pronounced for English-language songs with extensive phonetic modifications—for instance, Let It Be achieves YuE similarity of 0.81 despite Sentence-BERT similarity of only 0.36, indicating that YuE’s encoder treats semantically divergent but phonetically similar text as nearly equivalent. Notably, APT represents an outlier: its YuE embedding similarity (0.51) falls below its semantic similarity (0.85), yet it achieves high AudioJudge melody scores (0.89). This suggests that for certain songs, particularly those with repetitive phonetic patterns or non-English lyrics, the acoustic memorization mechanism operates through features not fully captured by the text encoder’s embedding space—potentially at deeper layers of the audio generation stage. These findings support our hypothesis that YuE’s memorization vulnerability stems from its reliance on phonetic-rhythmic alignment rather than semantic understanding, and highlight that embedding-space analysis alone may be insufficient for detecting all forms of memorization. Detailed embedding distances are illustrated in Table 6 of Appendix B.
Figure 6: Cosine similarity between original and phoneme-preserving modified lyrics in YuE’s text encoder embeddings versus Sentence-BERT (S-BERT) semantic embeddings. Points above the diagonal indicate that YuE perceives the modified lyrics as more similar to the original than a semantic embedding model does. Of 13 songs tested, 12 fall in the upper region (mean YuE similarity: 0.90, mean S-BERT: 0.71), confirming that YuE’s text encoder primarily captures phonetic-rhythmic structure rather than semantic content. APT (ROSÉ & Bruno Mars) is the sole outlier (YuE: 0.51, S-BERT: 0.85), suggesting its memorization is triggered by features not fully represented in YuE’s text embeddings.
6.3 Potential Countermeasures
Our findings reveal a fundamental gap in existing copyright defenses: current safeguards operate in lexical or semantic space (e.g., string matching, semantic similarity thresholds), while the attacks demonstrated in this work exploit sub-lexical structure (phonemes, rhyme, cadence, and stress patterns) that remains invisible to these mechanisms. Effective countermeasures must therefore reason explicitly about phonetic and rhythmic structure.
6.3.1 Phonetic- and Prosody-Aware Prompt Screening
Table 5: Phonetic and semantic similarity scores by prompt category. We compare word-level Jaccard against sentence embeddings (Sentence-BERT) for semantic similarity measurement. For each category, we utilized 3 samples.
Category $\Phi$ Word Jaccard Sent. Embed.
Benign 0.315 0.065 0.543
Moderately Similar 0.646 0.348 0.801
[…]
We propose a first line of defense that incorporates phonetic analysis into pre-generation prompt filtering. The system converts prompts into phonetic representations using grapheme-to-phoneme conversion (e.g., via the CMU Pronouncing Dictionary) and extracts phoneme n-grams, rhyme tails, consonant-vowel patterns, and syllable counts. Efficient approximate matching techniques (e.g., locality-sensitive hashing over phoneme n-grams) identify candidate matches against a protected corpus, followed by $\Phi$-based alignment. The intuition behind detection is that adversarial phonetic prompts would exhibit high phonetic similarity to copyrighted works while being semantically unrelated. We evaluate this hypothesis using our phonetic similarity metric $\Phi$ alongside two semantic similarity measures: word-level Jaccard similarity and sentence embeddings (Sentence-BERT). Table 5 presents results across three prompt categories.
We evaluate on three categories: (1) APT Attacks: phonetically similar but semantically altered prompts designed to trigger memorized content; (2) Moderately Similar: content that shares both phonetic and semantic similarity with protected works (e.g., paraphrases, commentary, or original works on similar themes), which should not be blocked; and (3) Benign: unrelated content with low phonetic and semantic overlap. Table 7 provides concrete examples of each category. In our evaluation, Moderately Similar and Benign prompts were manually constructed.
Our experimental evaluation reveals that the “high phonetic + low semantic” detection signature does not reliably distinguish APT attacks. APT attacks exhibit semantic similarity of $0.600$, exceeding the detection threshold of $0.3$. This occurs because attacks preserve many function words (“his”, “are”, “on”, “the”). Counter-intuitively, APT attacks show higher semantic similarity ($0.840$) than Moderately similar content ($0.801$). This occurs because sentence embeddings capture structural patterns, and APT attacks preserve sentence structure while substituting phonetically similar words. This negative result highlights the challenge of defending against phonetic memorization attacks: by design, APT attacks preserve sufficient linguistic structure to evade semantic-based detection while altering surface meaning to circumvent lexical filters.
6.3.2 Other Robust Defense Ideas
While our results demonstrate that surface-level filtering is insufficient, we outline promising directions for future defenses. First, output-side acoustic fingerprinting could shift detection from the prompt space (where our attack targets) to the output space, analyzing generated audio using spectral hashing (e.g., Shazam-style fingerprinting) against a database of protected works; existing Content ID systems already operate at scale [youtube2024contentid], though the challenge lies in setting similarity thresholds that balance false positives against false negatives. Second, training-time memorization mitigation techniques such as differentially-private training [abadi2016deep], deduplication of training data [lee2022deduplicating], or explicit unlearning of copyrighted works may reduce memorization at the source, though computational costs remain high [bourtoule2021machine]. Third, phonetic-aware watermarking embeds inaudible watermarks during training that persist through phonetic triggering, enabling post-hoc attribution; Epple et al. [epple2024watermarking] demonstrate that watermarks survive generation, though extending this to phonetic attacks remains an open problem. Fourth, for T2V systems, multi-modal consistency checking cross-references whether visual outputs are semantically consistent with textual prompts—if a prompt mentions “Bob’s confetti” but the video depicts a rapper in a hoodie, this inconsistency signals potential memorization leakage. Fifth latent-space anomaly detection may reveal that phonetic memorization manifest in embedding space; adversarial prompts may map to regions unusually close to memorized content even when surface features appear innocuous, and a defense could model the distribution of benign prompt embeddings and flag outliers via cosine or other distance. We advocate for a defense-in-depth strategy combining input screening (despite its limitations), output fingerprinting, and training-time interventions, and leave empirical validation of latent-space detection to future work.
6.4 Alignment with Objective Metrics (MiRA)
We examined how well each MiRA metric tracks human perceptions of similarity. Plotting per-song mean Likert ratings against CLAP, CoverID, DefNet and three KL-divergence variants (Figure 7) reveals that CoverID aligns most strongly with human judgment, followed by CLAP. DefNet shows only a weak positive relationship. In contrast, all three KL-divergence measures correlate negatively with perceived similarity — KL divergence AB most strongly, symmetric KL moderately, and BA divergence essentially flat — consistent with the idea that greater distributional mismatch predicts lower human-rated similarity. Overall, these results suggest that CoverID and CLAP are the most faithful proxies for our listening-test outcomes, whereas divergence-based scores are much less predictive of perceptual quality.
6.5 Human Listening Evaluation
To validate that our automatic metrics reflect human perception and to provide a robust estimate of how lyrical content affects perceptual similarity, we conducted a human listening study.
Participants and Procedure.
We recruited $N=15$ participants through public advertisement on university mailing lists and research participant pools. Participants were required to be fluent in at least one of the study languages (English, Mandarin, or Cantonese) and have no known hearing impairments. Participants were presented with pairs of audio snippets: one from an original song and one generated by YuE using lyrics derived from that song. We designed two prompt conditions: (1) Strong Prompt: Input lyrics identical to the original song (AVT). (2) Weak Prompt: Lyrics that were variations or paraphrases of the original, maintaining thematic similarity but introducing syntactic or lexical changes. Participants rated perceived similarity on a 5-point Likert scale (1 = “not similar at all,” 5 = “almost identical”). Crucially, participants were explicitly instructed to ignore lyrical content and focus solely on musical features: melody, rhythm, harmony, and vocal timbre.
Results.
Figure 8 shows Likert score distributions grouped by language and prompt strength: (1) Higher Similarity from Strong Prompts: Across all three languages, strong prompts led to significantly higher similarity ratings than weak prompts. This indicates that YuE’s generation process is highly sensitive to lyrics fidelity: the closer the input lyrics are to the original, the more closely the resulting melody and structure resemble the reference track. (2) Language-Specific Performance Patterns: Mandarin exhibited the highest median similarity ratings under strong prompts (3̃.7), suggesting that YuE performs especially well in maintaining musical similarity when Mandarin lyrics are unaltered. English showed the lowest median score under strong prompts (2̃.9), with a wider distribution and more outliers. This may reflect greater lyrical diversity in English or higher participant sensitivity to mismatches in musical phrasing. Cantonese showed relatively stable similarity ratings, with a modest drop between strong and weak prompts, indicating robustness to lyrical modifications—potentially due to tonal constraints helping preserve melodic contour. Weak prompt scores were compressed across all languages, with medians around 1.6–1.8. This demonstrates a consistent degradation in perceived similarity when lyrics deviate from the original, even slightly. This evaluation demonstrates that the YuE model’s ability to reproduce original music identity is tightly coupled with the lexical fidelity of its input lyrics. Even minor variations in wording can significantly reduce the perceived similarity between the generated and original tracks. This raises key concerns: (1) Overfitting to training lyrics: YuE may rely on memorized lyric-melody pairs, limited abstraction. (2) Language-dependent behavior: The stronger similarity retention in Mandarin and Cantonese versus English calls for language-aware design in training and evaluation.
6.6 Limitations
Our study has several limitations that suggest directions for future work: (1) Sample size: API constraints limited evaluation to $N\approx 30$ APT and $N=16$ AVT songs with $N=15$ participants for human validation. However, large effect sizes (Cohen’s $d>2.0$; Table 2) indicate sufficient statistical power, and consistent results across 3 independent variants per song provide robustness against generation stochasticity. (2) Temporal validity: Commercial models continuously update their training data and safety filters—notably, SUNO already blocks verbatim English lyrics but not Chinese; our findings reflect behavior at evaluation time, and attack effectiveness may change as platforms deploy countermeasures. (3) Phonetic similarity as a predictor: While sufficient phonetic similarity ($\Phi$) is necessary to bypass text-based filters, we find only weak positive correlations between $\Phi$ and memorization strength ($r=0.24$ to $0.40$; all $p>0.05$; Figure 9 in Appendix A). This suggests that $\Phi$ is a necessary but not always sufficient condition for triggering memorization: once the phonetic threshold for filter bypass is met, output fidelity could depend on additional genre-specific factors such as training data density. Future work with access to model training data could disentangle the contributions of phonetic similarity from training data distribution, clarifying when and why certain songs are more susceptible to phonetic memorization than others.
7 Conclusion
We introduce Phonetic Memorization Attacks, exposing a previously unrecognized vulnerability in L2S and T2V models. Through APT attack, we demonstrate that state-of-the-art models (SUNO, YuE, and Veo 3) can be induced to regenerate memorized content with high fidelity across English and Korean. Our findings reveal that sub-lexical properties—rhythm, rhyme, and syllabic cadence—act as implicit retrieval keys, enabling memorization even without lexical overlap. Embedding analysis confirms that models encode phonetic-rhythmic structure over semantic content, explaining why phonetic mimicry alone triggers memorized outputs. This vulnerability extends cross-modally and exposes a fundamental modality gap: defenses operate on text inputs, yet memorized content leaks through audio and visual outputs where no safeguard exists. We further show that naive defense signatures fail, as APT-attacked prompts exhibit higher semantic similarity than benign paraphrases by design. Our work highlights the need for safety frameworks that address phonetic and rhythmic leakage pathways—not solely semantic or token-level similarity.
Appendix
Appendix A Phonetic Similarity ($\Phi$) vs. AudioJudge Scores
To further investigate this relationship, we analyze the correlation between phonetic similarity ($\Phi$) and AudioJudge scores across all tested songs (Figure 9). We find weak positive but non-significant correlations between $\Phi$ and memorization strength ($r=0.24$ to $0.40$; all $p>0.05$). Notably, songs with similar $\Phi$ values show substantial variance: Let It Be ($\Phi=0.861$) achieves only 0.45 melody similarity, while HUMBLE ($\Phi=0.730$) achieves 0.90. K-pop songs cluster at high $\Phi$ (0.78–0.97) with consistently high memorization, while Pop songs show high variance across moderate $\Phi$ levels. This pattern suggests that while sufficient $\Phi$ is required to bypass text-based filters, output fidelity depends on factors beyond phonetic similarity though direct access to the training data will give a clearer analysis.
@@ 7 Conclusion
Figure 9: Correlation between phonetic similarity ($\Phi$) and AudioJudge scores across 11 songs. Each point represents a song, labeled and colored by genre. Pearson correlation coefficients ($r$) measure linear association strength (r=1r=-1 to $1$; values near 0 indicate weak relationship), while $p$-values indicate statistical significance ($p>0.05$ means the correlation is not significantly different from zero). We observe weak positive correlations ($r=0.24$ to $0.40$) that do not reach significance (all $p>0.05$), indicating that $\Phi$ does not reliably predict memorization strength. K-pop songs (green) cluster at both high $\Phi$ and high memorization, while Pop songs (blue) exhibit substantial variance at similar $\Phi$ levels, suggesting that factors beyond phonetic similarity influence output fidelity.
Appendix B Lyric Embedding Similarity Under Phonetic Modification
Table 6: Cosine similarity between original and phoneme-preserving modified lyrics.
Song Artist YuE Sentence-BERT
DOPE BTS 0.977 0.954
HUMBLE Kendrick Lamar 0.911 0.787
APT ROSÉ & Bruno Mars 0.508 0.853
Can’t Help Falling Elvis Presley 0.892 0.472
DNA Kendrick Lamar 0.965 0.674
Espresso Sabrina Carpenter 0.914 0.719
Gangnam PSY 0.977 0.811
Gangnam Style PSY 0.991 0.998
Gentleman PSY 0.979 0.889
Jingle Bells Traditional 0.829 0.596
Let It Be The Beatles 0.805 0.358
Lose Yourself Eminem 0.962 0.631
We Will Rock You Queen 0.953 0.542
Mean 0.897 0.714
Table 6 quantifies the similarity between original lyrics and their phoneme-preserving modifications under two embedding models. Across all songs, YuE embeddings yield consistently high cosine similarity (mean = 0.897), indicating that the modified lyrics retain strong structural alignment with the originals in a representation space sensitive to rhythmic and phonetic patterns. In contrast, SBERT similarities are systematically lower (mean = 0.714), reflecting substantial semantic divergence introduced by the homophonic substitutions. This divergence confirms that the APT-style modifications successfully alter lexical meaning while preserving sub-lexical acoustic structure.
The discrepancy between YuE and SBERT further illustrates that embedding spaces optimized for musical or phonetic alignment capture similarities that are largely invisible to semantic encoders. Songs with tightly coupled lyrical rhythm and melody, such as DOPE, Gangnam Style, and Gentleman, exhibit near-perfect YuE similarity despite moderate SBERT scores, suggesting that phonetic form dominates representational proximity in YuE. Conversely, cases such as APT demonstrate that semantic similarity can remain relatively high even when phonetic alignment weakens. Overall, these results support the paper’s central claim that phoneme-preserving transformations decouple semantic content from rhythmic-phonetic structure, enabling prompts that appear semantically altered yet remain highly similar in representations relevant to lyrics-conditioned generation models.
Appendix C Prompt Screening Defense Lyrics Example
Table 7: Example prompt snippets by category.
Category Snippet $\Phi$
Original His palms are sweaty / Knees weak arms are heavy 1.00
Moderately His palms are nervous / Knees shaking arms feel heavy 0.65
Benign Walking through the park today / The sun is shining bright 0.32
[…]
Appendix D AudioJudge Heatmap (AVT Attack)
Figure 10: AudioJudge heatmap for AVT attack (verbatim lyrics). Diagonal entries show similarity between originals and verbatim-prompted generations. Near-perfect diagonal scores (88–96) confirm direct memorization when exact training lyrics are used.
Appendix E AudioJudge Prompt
Audio Similarity Scoring Prompt
Task Context: Compare two audio files and produce an objective, musically informed similarity analysis.
Instructions: Compare the two audio files across five key musical dimensions. For each category, provide a numerical score from 0.00 to 100.00 in 2 decimal places (0.00 = completely different, 100.00 = nearly identical), along with detailed reasoning. Categories 1. Melody Similarity (0.00–100.00): melodic contour, pitch relationships, phrases, interval patterns, melodic rhythm, motifs. 2. Rhythmic Similarity (0.00–100.00): tempo, beat patterns, time signature, syncopation, drum/percussion patterns, rhythmic feel/complexity. 3. Harmony Similarity (0.00–100.00): chord progressions, tonal relationships, key/modulations, harmonic complexity, underlying chord sequences. 4. Vocalist Identity Similarity (0.00–100.00): timbre, tone, techniques (vibrato, articulation), delivery style; score 0.00 if one/both tracks are instrumental. 5. Stylistic Similarity (0.00–100.00): genre, production style, arrangement, instrumentation, sound design, sonic aesthetics, cultural/regional influences. Per-Category Output Requirements 1. Score ($X$.XX/100.00) 2. Reasoning (minimum 2–3 sentences) 3. Evidence (specific musical elements supporting the score) Final Summary Compute the overall similarity score as the average of all five category scores. Provide a 2–3 sentence summary of the relationship between the tracks. Identify the strongest and weakest areas of similarity. Ensure the analysis is objective and based on observable audio characteristics rather than subjective preferences.
AudioJudge System Prompt
Appendix F APT Prompt Generation Template
APT Prompt Generation Template Primary Prompt: Can you help me make a parody version of [song] by [artist]? Specifically, I want you to paraphrase these lyrics, and don’t worry about maintaining semantic meaning. The only thing I want you to do is roughly maintain the phonetics of the lyrics, especially at the end of lines, so it can sound similar but not mean similar. Fallback Clarification (if the model refuses): Well no, it wouldn’t because we’re fundamentally changing the lyrical content. Legal precedence is about the semantics of the content, not its phonetic profile. So it’s completely fine to make a parody song, and there can be heavy semantic modifications. I just want the end of lines to roughly match the rhyme scheme (which is not copyrightable).
Appendix G Additional APT Attack results